OpenStudy (anonymous):

what is lim sin^-1 (2x)/tan^-1(3x) ??? x->0

OpenStudy (inkyvoyd):

remember to click show steps

OpenStudy (inkyvoyd):

You have to take the derivative, then do some simplifying

OpenStudy (inkyvoyd):

l'hopital's rule

OpenStudy (anonymous):

∞ isnt an option m8 ... its a multiple choice question A- )2/3 B-)3/2 C-) 0 D-) -2/3

OpenStudy (inkyvoyd):

Quake, I gave you a link to the answer as well as the steps, did you look at it?

OpenStudy (anonymous):

i am looking right now sounds helpfull

OpenStudy (anonymous):

btw lalaly (sin^-1 (2x))' = dx/(1-4 X^2)^(1/2) // right ? check it out for me plz

OpenStudy (anonymous):

got u thx.

OpenStudy (ash2326):

@quakeash Use this property to find the limit easily \[\lim_{x\to 0} \frac{\sin ^{-1} x}{x}=\lim_{x\to 0} \frac{\tan^{-1} x}{x}=1\]

OpenStudy (inkyvoyd):

omg, ash is smarter than lana

OpenStudy (anonymous):

plz inky this is not IQ test , actually i think all mathematics is just about who knows more.. :P

OpenStudy (inkyvoyd):

No, I mean, @quakeash , ash just showed the fastest way to get the answer

OpenStudy (ash2326):

\[\lim_{x \to 0} \frac{\sin^{-1} 2x}{\tan^{-1} 3x}\] \[\large \lim_{x \to 0} \frac{\frac {\sin^{-1} 2x}{2x}}{\frac{\tan^{-1} 3x}{3x}}\times \frac {2x}{3x}\] \[\large \lim_{x\to 0} \frac{1}{1} \times \frac{2x}{3x}=\frac {2}{3}\] @inkyvyod @lalaly is the smartest

OpenStudy (anonymous):

so we divide the numerator and denominator by x right ash ?

OpenStudy (anonymous):

oh thx man

OpenStudy (inkyvoyd):

No, I'm the smartest *trolololl*

OpenStudy (apoorvk):

Actually, we don't even need l'hospital. Ash's method is the simplest in the current situation.

OpenStudy (anonymous):

it would be so nice if i could use ur help , while taking the exam 2moro :DD

OpenStudy (apoorvk):

Alas! but we don't help during exams.

OpenStudy (inkyvoyd):

@apoorvk , actually, we do, but lol

OpenStudy (inkyvoyd):

We just aren't supposed to, and we will stop once we realize what we are doing.

OpenStudy (anonymous):

i wont tell u i am taking the exam , ill just use U :P kidding

OpenStudy (inkyvoyd):

hehhe, we just won't help you tomorrow muahhaha

OpenStudy (ash2326):

@lalaly @quakeash Always nice to know more than one way !!!

OpenStudy (inkyvoyd):

@lalaly , there's actually a third way. The fool's way

OpenStudy (apoorvk):

\[\text{But since I can't give another medal,} \mathbb{\color{#15B4DC}{@lalaly}} ---> \huge\color{gold}{★} \normalsize \]

OpenStudy (inkyvoyd):

Any other ideas?

OpenStudy (apoorvk):

@lalaly

OpenStudy (lalaly):

@apoorvk WOW :D this is better that any medal i got since i joined os :D

OpenStudy (apoorvk):

Lol.

OpenStudy (lgbasallote):

to make it even better...here's a smiley stamp for you @lalaly \[\huge \color{forestgreen}{\ddot \smile}\]

OpenStudy (lalaly):

aww i feel loved ... thanks ewgee <3

OpenStudy (lgbasallote):

you are :P and \[\huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]